For the following reaction, find the moles of product made given 5.41 g O2 and 2.71 g of Al. Use "mole" as unit. Do NOT use scientific notation in answer.

Al + O2 --> Al2O3

Atomic mass: Al = 26.98 g/mol, O = 16.00 g/mol


1
Expert's answer
2021-06-23T05:16:55-0400

Molar Mass of Al = 26.98

Ratio = 1:2

2.71 = 2.71/26.98

= 0.1 mole × 2 = 0.2mole

Molar Mass of O = 16

Ratio = 2:3

= 5.41/16

= 0.338 × 3/2

= 0.51 mole

= 0.51 + 0.2

= 0.71 mole